# Generative User-Experience Research for Developing Domain-specific Natural Language Processing Applications

### v1

@@ 8 Conclusion
The paper proposes a methodology for integrating generative UX research into a process of developing domain-specific NLP applications. Generative UX research employs domain users at the initial stages of prototype development, i.e., ideation and concept evaluation, and the last stage of the evaluation. The core idea of generative research is that involving domain users increases their interest and trust in the final application since the application addresses their needs.
As a case study, we report a full cycle of the prototype development of a domain-specific NLP application for daily operations in the process industry. In the generative UX research, we learn about a complicated domain German language, varying writing styles, partial data incompleteness in the logging system, and the short lifetime of the text records. We propose a semantic search to enable effective access to the historical knowledge of plant operations and motivate users to improve data completeness by logging more information. The user-oriented prototype evaluation focuses on assessing the accuracy and user value after testing the prototype. The prototype evaluation demonstrates more informed, effective, and efficient daily operations.
To sum up, the paper shows that generative research for NLP application development is a combination of the strengths of both UX and NLP research. The synergy reveals opportunities and constraints of the data- and user-driven nature through exploratory data analysis and contextual inquiries. Working in close collaboration, our NLP+UX team and the domain experts created and validated a concept idea for a prototype through combined prototype evaluation of accuracy and user value.

### v4

@@ 6 Conclusion
The paper reports on a case study of developing a domain-specific NLP application prototype in the process industry. To develop a prototype, we proposed a methodology for integrating generative UX research into the data-driven methodology of developing NLP applications by reinforcing domain understanding through data analysis and user interviews. Generative UX research focuses on domain users and employs them at the initial stages of prototype development, i.e., ideation and concept evaluation and the last stage of the prototype evaluation. The core idea of generative research is that involving domain users increases user utility, interest, and trust in the final application since it targets their needs.
The case study reports a full prototype development cycle of a domain-specific NLP application for daily operations in the process industry. In the generative UX research, we learn about a complicated domain of the German language, varying writing styles, partial data incompleteness in the logging system, and the short lifetime of the text records. We propose a semantic search to enable effective access to the historical knowledge of plant operations and motivate users to improve data completeness by logging more information. The user-oriented prototype evaluation focuses on assessing the accuracy and user utility when interacting with the prototype with a set of provided tasks or openly as they see fit. The prototype evaluation demonstrates more informed, effective, and efficient task completion in daily operations.
The paper shows that generative research for NLP application development combines the strengths of both UX and NLP research. The methodology is a promising direction for broad use in the development of NLP applications. The UX+NLP combination reveals opportunities and constraints of the data- and user-driven nature through exploratory data analysis and contextual inquiries, which enables a more informed and effective development process. We encourage UX+NLP teams to work in close collaboration to test the proposed methodology and report their results.